A girl weighing 470N hangs suspended on the end of a rope 8 meters long. What will be her gain in potential energy when a boy swings her to one side so that the rope makes an angle of 35º with the vertical?


1
Expert's answer
2022-03-03T14:59:28-0500
"\Delta E_p=mg\Delta h=mgl(1-\cos\theta)"

"\Delta E_p=\rm 470\: N*8\: m*(1-\cos35^\circ)=680\: J"